### v1

@@ II-C Smart Walking Aids and Applied Sensor Technologies
Walking aids play a crucial role in daily mobility support for patients with walking impairments. Recent research has explored how these devices can be enhanced through sensor technologies to support health monitoring, user interaction, and feedback-based interventions [74, 75]. Various smart walking aid prototypes have been proposed, including instrumented sticks [76], canes [77], and forearm crutches [78]. Many of these systems have been developed specifically for users with visual impairments [79, 80, 81, 82]. These systems typically contain sensors such as IMUs for fall detection or for motion tracking [83, 84], FSRs embedded in handles to detect gait events [85], GPS modules for positional tracking [86], and ultrasonic sensors for obstacle detection [79, 80, 81].
@@
Beyond data acquisition, several studies have explored the integration of feedback mechanisms to enhance user support [82, 87, 88]. For example, Merrett et al. [87] demonstrated the potential of biofeedback in forearm crutches using acoustic signals based on FSR data to support optimal weight distribution. Schuster et al. [88] showed that haptic feedback delivered via the handle of a cane can reduce knee adduction moments in individuals with knee osteoarthritis. In addition, a smart stick with integrated SOS functionality has been developed that connects to smartphone applications [89].
@@
While individual sensor-based walking aids have demonstrated promising results, integrated systems that combine smart walking aids with intelligent footwear for foot augmentation remain rare and are typically limited to specific application scenarios. For instance, Suman et al. [90] proposed a system in which a smart cane communicates with smart footwear, using ultrasonic sensors in the shoe and auditory feedback via the cane handle. Another example is the use of force-sensitive crutches to assist lower-limb exoskeleton users by capturing ground force data to enable active support [91]. Furthermore, a robotic walking cane was developed that receives center of pressure (COP) data from four FSRs integrated in an insole to support fall detection [83]. However, direct communication between walking aids and orthopedic devices such as sensor-based foot orthoses has not yet been realized. In this context, Resch et al. [14] proposed a conceptual system that enables a walking aid to communicate with a smart AFO through vibration feedback in the handle to indicate plantar pressure overload. Nevertheless, the proposed system remains a concept without implementation or empirical evaluation.
@@
Overall, while sensor-based walking aids and feedback modalities offer high potential, most systems remain in the prototype stage and lack systematic validation or integration into broader rehabilitation frameworks [75].
II-D Research Gap
@@
Related work lacks concepts and empirical evidence for the combined use of smart walking aids and foot monitoring technologies. However, this integration offers future potential for improving patient monitoring, supporting real-time gait analysis and enabling multimodal feedback in healthcare and everyday use. Building on these findings, our work addresses these gaps by developing and validating an integrated system that combines a smart foot orthosis, an instrumented walking aid, and a smartphone application for real-time gait monitoring and interactive feedback. By addressing these gaps we contribute to the advancement of assistive device technologies, mobile gait analysis, and user-centered feedback systems.
III Method
This section describes the overall system architecture and the technical implementation of the smart orthosis, the instrumented walking aid, and the smartphone application. It is structured as follows: Section III-A provides a system overview, Section III-B presents the hardware architecture and concept development, and Section III-C outlines the software architecture and data processing for each subsystem. Finally, Section III-D presents the experimental setup of the prototype validation study, including apparatus, procedure, and measurement with data analysis.
III-A System Overview
@@
The SFO is based on a standardized lower-leg orthosis (AIRCAST® AIRSELECT™ Elite Walker  [^1] https://www.enovis-medtech.de/AIRCAST-AIRSELECT-Elite-Walker by ORMED GmbH), also known as a controlled ankle motion boot, which can be worn universally on either foot. A sensor concept consisting of multiple FSRs and a single IMU is integrated into a 3D-printed insole to capture plantar pressure and motion data. The insole is designed as a modular component, with adjustable size and shape to accommodate different orthotic devices and foot sizes.
Sensor fusion and real-time data processing enable the extraction of gait parameters and the detection of critical loading patterns. To support active user guidance, haptic feedback is provided via a vibration motor embedded in the orthosis, which is triggered in response to excessive plantar pressure. A custom-designed printed circuit board (PCB), equipped with an on-board microcontroller unit (MCU), enables wireless communication via Bluetooth Low Energy (BLE) with both the mobile application and the instrumented walking aid. The forearm crutch is based on a standardized design (ISO 11334-1:2007) and includes a custom-designed PCB equipped with an MCU, an IMU sensor, two vibration modules, and a rechargeable battery. All components are embedded in the handle to enable synchronized haptic feedback and motion sensing. In addition, a corresponding mHealth application developed in Android Studio provides a graphical interface for visualizing gait parameters and plantar pressure data.
III-B Hardware Architecture and Concept Development
III-B1 Smart Orthosis
A schematic overview of the SFO prototype is presented in Figure 2(a) and the assembled insole with integrated sensor components is shown in Figure 2(b).
(a) Overview of the orthosis prototype and assembled components.
(b) Final assembled prototype of the smart insole.
Figure 2: Schematic and physical implementation of the smart foot orthosis.
@@ III-B1 Smart Orthosis
The SFO is based on a modular insole integrated into a standard lower-leg orthosis. An optional external vibration module can be placed at varying heights along the calf using the holes of the orthosis, allowing flexible positioning for haptic feedback. The insole includes multiple FSRs and an IMU sensor, connected via a flat cable to an external PCB, which is enclosed in a 3D-printed housing mounted on the posterior strap of the orthosis. The enclosure was designed using the computer-aided design (CAD) software Autodesk Fusion 360 and fabricated from polylactid (PLA) with a 0.4 mm nozzle (HT extruder) using the Flashforge Creator 4 3D printer. The insole consists of a two-part structure, comprising a bottom and a top layer, both fabricated from flexible thermoplastic polyurethane (TPU) using the extruder F with 0.4 mm nozzle size. The geometry replicates the original insole of the AIRCAST® orthosis to ensure proper fit and alignment. The design enables adaptation for the contralateral healthy foot by modifying the insole shape and enclosure mounting to fit standard footwear. It also supports future integration into other orthotic systems tailored to individual foot anatomies.
@@
The modular insole was designed to include the recommended minimum of 15 FSRs [42] (Model 400 short tail, Interlink Electronics [^2] https://www.interlinkelectronics.com/fsr-400-short) to capture plantar pressure distribution. The placement of the FSRs is based on the anatomical structure of the foot and covers key plantar regions, including the heel, forefoot, metatarsal arch, and toes. The layout supports extensions up to 24 FSRs, which may be beneficial for applications involving foot deformities or more detailed regional pressure mapping. Each sensor has an active area of 5.6 mm in diameter, a thickness of 0.3 mm, and a force sensitivity range of 0–20 N, offering fast response times and high reliability for dynamic pressure measurements [92].
@@
For motion tracking, a 6-axis IMU (BMI270, Bosch [^3] https://www.bosch-sensortec.com/products/motion-sensors/imus/bmi270/) was integrated into the insole. The IMU was positioned near the metatarsal arch to minimize pressure exposure. This sensor was selected based on key technical requirements for wearable gait analysis systems and is commonly used in fitness trackers and smart wearables, such as smart shoes for step counting and activity recognition [93]. The ultra-low power sensor combines a 3-axis accelerometer (sampling frequencies from 25 Hz to 6.4 kHz) and a 3-axis gyroscope (0.78 Hz to 1.6 kHz) within compact dimensions of 0.8 mm $\times$ 3 mm $\times$ 2.5 mm.
@@
To minimize space, components were placed and routed according to system requirements and manufacturer guidelines. The final design was verified using electrical rule check (ERC) and design rule checks (DRC). Board assembly was performed manually using solder paste and a heating plate for surface-mounted device (SMD) soldering, followed by manual mounting of through-hole technology (THT) components. The assembled board was visually inspected and functionally tested to ensure proper soldering and reliable system performance. Additionally, detailed battery tests confirmed the functionality of the integrated battery management system, including discharge and charging profiles. The estimated runtime without vibration feedback was approximately 35.5 hours, with a full recharge time of 6–7 hours depending on the initial state of discharge.
III-B2 Smart Walking Aid
(a) Hardware block diagram of the smart walking aid.
(b) Final assembled prototype with integrated electronics.
Figure 3: Schematic and physical implementation of the instrumented walking aid.
To enable active haptic feedback through an assistive device, a forearm crutch was selected as the basis for sensor integration, as such walking aids are commonly used in combination with AFOs and offer a practical interface for feedback delivery. We developed a modular integration concept that can be transferred to various walking aids, including sticks, canes, or walkers, provided the handle design allows for modification without compromising ergonomics. This is essential, as major modifications to the handgrip require careful consideration of weight distribution and ergonomic design to avoid negative impact. For example, sensor placement along the shaft could alter the center of gravity, which may require new validation/certification of the walking aid [94]. Therefore, our design specifically targeted integration without major modifications to the handgrip. In this implementation, we used a standardized aluminum forearm crutch with a plastic grip and an integrated reflector. After removing the reflector, the inner cavity of the handle provided a usable space of 25 mm in diameter and 115 mm in length for the integration of the electronic components.
@@ III-B2 Smart Walking Aid
The USB connector, MCU, and battery holder were symmetrically arranged along the central axis of the PCB to ensure balanced layout and optimal use of space. The MCU and USB connector were placed on the front side, while the battery was positioned on the back side in accordance with the manufacturer’s mounting guidelines. A universal asynchronous receiver transmitter (UART) interface was included to allow communication with an external PC and GUI during development. Before fabrication, ERC and DRC were performed to verify the circuit design and layout. The PCB was then assembled using a manual pick-and-place process, followed by reflow soldering in a three-zone convection oven (Mistral 260) for SMD components. Additional THT components were soldered manually in a subsequent step. After assembly, the boards underwent functional testing and visual inspection to ensure proper soldering and system performance. To ensure optimal positioning and mechanical stability within the handle, the assembled PCB was mounted inside a custom 3D-printed enclosure made of rigid PLA. To enable charging without disassembly, a borehole was added to the back of the crutch, providing external access to the USB-C port. An overview of the hardware architecture is provided in Figure 3(a), and the assembled prototype is shown in Figure 3(b).
III-C Software Architecture and Data Processing
III-C1 Smart Orthosis - Mobile Gait Analysis
Figure 4 presents the SFO data processing architecture, structured into five stages: (1) raw data acquisition, (2) pre-processing, (3) signal processing, (4) sensor fusion, and (5) extraction of gait parameters. Each step is described in detail in the following paragraphs.
@@ III-C1 Smart Orthosis - Mobile Gait Analysis
Raw Data Acquisition
[…]
Pre-processing
[…]
$V_{\text{out}}=\frac{R_{M}\cdot V^{+}}{R_{M}+R_{\text{FSR}}}$ (1)
Following the manufacturer’s specification [92], the raw sensor output was used to estimate relative plantar forces without applying sensor-specific calibration. To reduce high-frequency noise, a moving average filter was applied to the estimated force signals.
The BMI270 includes internal configuration options for filtering and calibration, which were utilized to enhance signal quality. The sensor was configured with a measurement range of $\pm 4g$ for the acceleration and $\pm 2000^{\circ}/s$ for the gyroscope to capture typical gait-related movements as well as rapid rotational changes. Offset correction was automatically performed during initialization by fixing the IMU in a stationary position to determine the static bias for each axis. These offset values were removed from all subsequent accelerometer and gyroscope samples to ensure base-corrected measurements. Additionally, separate internal low-pass filters were applied to the accelerometer and gyroscope axis to remove high-frequency noise. Pre-processing of the IMU data through offset correction and low-pass filtering ensures that the acceleration and rotation data are precise and reliable.
Signal Processing
In the signal processing stage, pre-processed sensor data were further refined to reduce residual noise and enable robust detection of gait events. To improve signal quality, Kalman filters [98] were applied to both FSR and IMU signals. The Kalman filter was selected due to its proven suitability for real-time gait analysis in wearable systems, including IMU-based orientation estimation, walking distance calculation, and general gait assessment [54, 99, 100]. For the FSR signals, individual Kalman filters were assigned to each of the sensor channels. The filter parameters were adjusted to maintain responsiveness while reducing high-frequency fluctuations. For the IMU, each axis of the accelerometer and gyroscope data was filtered individually, with configuration parameters adjusted to the sensor’s noise characteristics.
The filtered FSR signals served as the basis for gait event detection. A custom rule-based algorithm was implemented to identify key gait phases using threshold values defined as relative percentages of the user’s body weight. This individualized approach enabled the detection of characteristic loading patterns and supported the identification of critical gait events and potential overload conditions. Thresholds were empirically determined through pre-testing across multiple gait cycles and optimized to minimize detection errors. A finite state machine was implemented to manage gait phase transitions based on these events. This approach ensured a consistent event sequence and enabled accurate temporal segmentation. A baseline correction algorithm was used to counteract the hysteresis effects of the FSR sensors and to ensure zero-base alignment after the load is removed. This adjustment was essential for maintaining continuous and reliable gait event detection across steps. The detected gait events, including their timestamps, were subsequently used to compute spatiotemporal gait parameters such as step duration, stance/swing phase duration, and cadence.
Sensor Fusion
Filtered FSR and IMU data were fused to extract both temporal and spatial characteristics of the gait cycle. While FSR data provide reliable information on foot-ground contact timing, the IMU captures the foot’s three-dimensional orientation and motion dynamics. A Mahony filter [101] was employed to estimate foot orientation in terms of roll, pitch, and yaw. This approach combines gyroscope data for short-term tracking with accelerometer data for long-term correction, effectively compensating for drift and high-frequency noise. To improve the accuracy of position estimation, zero-velocity updates [102] were applied during the stance phase, identified from FSR signals. These updates reset the velocity to zero between the heel contact and the heel off, thereby limiting integration drift and reducing cumulative error in position calculation.
Extraction of Gait Parameters
Based on the fused sensor data, key gait parameters were extracted. The segmentation of the gait cycle follows the definition of Perry [103], which distinguishes two main phases: the stance phase (60% of the gait cycle) and the swing phase (40%), as well as eight sub-phases (including foot strike, foot flat, heel off, and foot off). These gait parameters can be classified into spatial, temporal, spatio-temporal, and kinetic categories [104]. Step length was estimated by integrating acceleration data, corrected for foot orientation and stabilized using zero-velocity updates during stance phases. Gait speed was subsequently derived as the ratio of step length to gait cycle duration. In addition to standard parameters, a stability index was introduced to quantify gait stability. This metric integrates two normalized components:
III-C2 Smart Walking Aid - Haptic Feedback and Motion Data

### v2

@@ 2.3 Smart Walking Aids and Applied Sensor Technologies
Walking aids play a crucial role in daily mobility support for patients with walking impairments. Recent research has explored how these devices can be enhanced through sensor technologies to support health monitoring, user interaction, and feedback-based interventions Yan et al. (2020); Resch et al. (2025). Various smart walking aid prototypes have been proposed, including instrumented sticks Srinivasan and Rajesh (2019), canes Yang et al. (2024), and forearm crutches Chamorro-Moriana et al. (2016). Many of these systems have been developed specifically for users with visual impairments Shah et al. (2021); Bansal et al. (2020); Ikbal et al. (2018); Gallo et al. (2010). These systems typically contain sensors such as IMUs for fall detection or for motion tracking Di et al. (2016); Tsuda et al. (2010), FSRs embedded in handles to detect gait events Trujillo-León et al. (2021), GPS modules for positional tracking Murali et al. (2017), and ultrasonic sensors for obstacle detection Shah et al. (2021); Bansal et al. (2020); Ikbal et al. (2018).
@@
Beyond data acquisition, several studies have explored the integration of feedback mechanisms to enhance user support Gallo et al. (2010); Merrett et al. (2010); Schuster et al. (2021). For example, Merrett et al. Merrett et al. (2010) demonstrated the potential of biofeedback in forearm crutches using acoustic signals based on FSR data to support optimal weight distribution. Schuster et al. Schuster et al. (2021) showed that haptic feedback delivered via the handle of a cane can reduce knee adduction moments in individuals with knee osteoarthritis. In addition, a smart stick with integrated SOS functionality has been developed that connects to smartphone applications Mohapatra et al. (2018).
@@
While individual sensor-based walking aids have demonstrated promising results, integrated systems that combine smart walking aids with intelligent footwear for foot augmentation remain rare and are typically limited to specific application scenarios. For instance, Suman et al. Suman et al. (2022) proposed a system in which a smart cane communicates with smart footwear, using ultrasonic sensors in the shoe and auditory feedback via the cane handle. Another example is the use of force-sensitive crutches to assist lower-limb exoskeleton users by capturing ground force data to enable active support Lancini et al. (2016). Furthermore, a robotic walking cane was developed that receives center of pressure (COP) data from four FSRs integrated in an insole to support fall detection Di et al. (2016). However, direct communication between walking aids and orthopedic devices, such as sensor-based foot orthoses, has not yet been realized. In this context, Resch et al. Resch et al. (2024) proposed a conceptual system that enables a walking aid to communicate with a smart AFO through vibration feedback in the handle to indicate plantar pressure overload. Nevertheless, the proposed system remains a concept without implementation or empirical evaluation.
@@
Overall, while sensor-based walking aids and feedback modalities offer high potential, most systems remain in the prototype stage and lack systematic validation or integration into broader rehabilitation frameworks Resch et al. (2025).
2.4 Research Gap
@@
Related work lacks concepts and empirical evidence for the combined use of smart walking aids and foot monitoring technologies. However, this integration offers future potential for improving patient monitoring, supporting real-time gait analysis and enabling multimodal feedback in healthcare and everyday use. Building on these findings, our work addresses these gaps by developing and validating an integrated system that combines a smart foot orthosis, an instrumented walking aid, and a smartphone application for real-time gait monitoring and interactive feedback. By addressing these gaps, we contribute to the advancement of assistive device technologies, mobile gait analysis, and user-centered feedback systems.
3 Method
This section describes the overall system architecture and the technical implementation of the smart orthosis, the instrumented walking aid, and the smartphone application. It is structured as follows: Section 3.1 provides a system overview, Section 3.2 presents the hardware architecture and concept development, and Section 3.3 outlines the software architecture and data processing for each subsystem. Finally, Section 3.4 presents the experimental setup of the prototype validation study, including apparatus, procedure, and measurements with data analysis.
3.1 System Overview
@@
The SFO is based on a standardized lower-leg orthosis (AIRCAST® AIRSELECT™ Elite Walker Enovis by ORMED GmbH, a company of Enovis, Freiburg, Germany), also known as a controlled ankle motion boot, which can be worn universally on either foot. A sensor concept consisting of multiple FSRs and a single IMU is integrated into a 3D-printed insole to capture plantar pressure and motion data. The insole is designed as a modular component, with adjustable size and shape to accommodate different orthotic devices and foot sizes.
Sensor fusion and real-time data processing enable the extraction of gait parameters and the detection of critical loading patterns. To support active user guidance, haptic feedback is provided via a vibration motor embedded in the orthosis, which is triggered in response to excessive plantar pressure. A custom-designed printed circuit board (PCB), equipped with an on-board microcontroller unit (MCU), enables wireless communication via Bluetooth Low Energy (BLE) with both the mobile application and the instrumented walking aid. The forearm crutch is based on a standardized design (ISO 11334-1:2007) NORM and includes a custom-designed PCB equipped with an MCU, an IMU sensor, two vibration modules, and a rechargeable battery. All components are embedded in the handle to enable synchronized haptic feedback and motion sensing. In addition, a corresponding mHealth application developed in Android Studio provides a graphical interface for visualizing gait parameters and plantar pressure data.
@@
3.2 Hardware Architecture and Concept Development
3.2.1 Smart Orthosis
@@
A schematic overview of the SFO prototype is presented in Figure 2a and the assembled insole with integrated sensor components is shown in Figure 2b. The SFO is based on a modular insole integrated into a standard lower-leg orthosis. An optional external vibration module can be placed at varying heights along the calf using the holes of the orthosis, allowing flexible positioning for haptic feedback. The insole includes multiple FSRs and an IMU sensor, connected via a flat cable to an external PCB, which is enclosed in a 3D-printed housing mounted on the posterior strap of the orthosis. The enclosure was designed using the computer-aided design (CAD) software Autodesk Fusion 360 (v.2603.0.86) and fabricated from polylactid (PLA) with a 0.4 mm nozzle (HT extruder) using the Flashforge Creator 4 3D printer. The insole consists of a two-part structure, comprising a bottom and a top layer, both fabricated from flexible thermoplastic polyurethane (TPU) using the extruder F with 0.4 mm nozzle size. The geometry replicates the original insole of the AIRCAST® orthosis to ensure proper fit and alignment, without affecting wearing comfort. The design enables adaptation for the contralateral healthy foot by modifying the insole shape and enclosure mounting to fit standard footwear. It also supports future integration into other orthotic systems tailored to individual foot anatomies.
@@
The modular insole was designed to include the recommended minimum of 15 FSRs Razak et al. (2012) (Model 400 short tail, Interlink Electronics Electronics ) to capture plantar pressure distribution. The placement of the FSRs is based on the anatomical structure of the foot and covers key plantar regions, including the heel, forefoot, metatarsal arch, and toes. The layout supports extensions up to 24 FSRs, which may be beneficial for applications involving foot deformities or more detailed regional pressure mapping. Each sensor has an active area of 5.6 mm in diameter, a thickness of 0.3 mm, and a force sensitivity range of 0–20 N, offering fast response times and high reliability for dynamic pressure measurements Interlink Electronics (2013).
((a))
((b))
Figure 2: (a) Overview of the orthosis prototype and assembled components; (b) Final assembled prototype of the smart insole.
@@
For motion tracking, a 6-axis IMU (BMI270, Bosch GmbH ) was integrated into the insole. The IMU was positioned near the metatarsal arch to minimize pressure exposure. This sensor was selected based on key technical requirements for wearable gait analysis systems and is commonly used in fitness trackers and smart wearables, such as smart shoes for step counting and activity recognition Bosch Sensortec GmbH (2023). The ultra-low power sensor combines a 3-axis accelerometer (sampling frequencies from 25 Hz to 6.4 kHz) and a 3-axis gyroscope (0.78 Hz to 1.6 kHz) within compact dimensions of 0.8 mm $\times$ 3 mm $\times$ 2.5 mm.
@@
To minimize space, components were placed and routed according to system requirements and manufacturer guidelines. The final design was verified using electrical rule check (ERC) and design rule checks (DRC). Board assembly was performed manually using solder paste and a heating plate for surface-mounted device (SMD) soldering, followed by manual mounting of through-hole technology (THT) components. The assembled board was visually inspected and functionally tested to ensure proper soldering and reliable system performance. Additionally, detailed battery tests confirmed the functionality of the integrated battery management system, including discharge and charging profiles. The estimated runtime without vibration feedback was approximately 35.5 h, with a full recharge time of 6–7 h depending on the initial state of discharge.
3.2.2 Smart Walking Aid
To enable active haptic feedback through an assistive device, a forearm crutch was selected as the basis for sensor integration, as such walking aids are commonly used in combination with AFOs and offer a practical interface for feedback delivery. We developed a modular integration concept that can be transferred to various walking aids, including sticks, canes, or walkers, provided the handle design allows for modification without compromising ergonomics. This is essential, as major modifications to the handgrip require careful consideration of weight distribution and ergonomic design to avoid negative impact. For example, sensor placement along the shaft could alter the center of gravity, which may require new validation/certification of the walking aid Ballesteros et al. (2019). Therefore, our design specifically targeted integration without major modifications to the handgrip. In this implementation, we used a standardized aluminum forearm crutch with a plastic grip and an integrated reflector. After removing the reflector, the inner cavity of the handle provided a usable space of 25 mm in diameter and 115 mm in length for the integration of the electronic components.
@@ 3.2.2 Smart Walking Aid
The USB connector, MCU, and battery holder were symmetrically arranged along the central axis of the PCB to ensure balanced layout and optimal use of space. The MCU and USB connector were placed on the front side, while the battery was positioned on the back side in accordance with the manufacturer’s mounting guidelines. A universal asynchronous receiver transmitter (UART) interface was included to allow communication with an external PC and GUI during development. Before fabrication, ERC and DRC were performed to verify the circuit design and layout. The PCB was then assembled using a manual pick-and-place process, followed by reflow soldering in a three-zone convection oven (Mistral 260) for SMD components. Additional THT components were soldered manually in a subsequent step. After assembly, the boards underwent functional testing and visual inspection to ensure proper soldering and system performance. To ensure optimal positioning and mechanical stability within the handle, the assembled PCB was mounted inside a custom 3D-printed enclosure made of rigid PLA. To enable charging without disassembly, an access was added to the back of the crutch, providing external access to the USB-C port. An overview of the hardware architecture is provided in Figure 3a, and the assembled prototype is shown in Figure 3b.
((a))
((b))
Figure 3: (a) Hardware block diagram of the smart walking aid; (b) Final assembled prototype with integrated electronics.
3.3 Software Architecture and Data Processing
3.3.1 Smart Orthosis—Mobile Gait Analysis
Figure 4 presents the SFO data processing architecture, structured into five stages: (1) raw data acquisition, (2) pre-processing, (3) signal processing, (4) sensor fusion, and (5) extraction of gait parameters. Each step is described in detail; see Appendix A.
@@ 3.3.1 Smart Orthosis—Mobile Gait Analysis
3.3.2 Smart Walking Aid—Haptic Feedback and Motion Data